# AARK: An Open Toolkit for Autonomous Racing Research

### v1

@@ V-A2 Vision
The datasets used to train ACMPC’s perception model are outlined in Table I. These provide semantic, depth and normal map annotations for each of AARK’s supported racetracks. Semantics consist of nine race track specific classes. Seven of these classes describe the type of surface: road, drivable, curb, carpet, grass, sand and water. Drivable is reserved to delineate the portion of road surface that is legal for vehicles to traverse. A track limits label is included that allows models to perceive road markings which can indicate different features of importance around a racecourse. The remaining classes denote physical objects in the environment: people, vehicles, structures and vegetation. Structures encompass a super set of buildings, fences, safety barriers and signs. ACDG is capable of generating ground truth annotations for the following machine vision tasks: semantic segmentation, depth, normal maps, semantic LiDAR, and 3D object detection. Currently AARK provides pre-generated datasets for the first three of these tasks.
@@ VI-A ACMPC
Often the barrier to entry into fields, such as autonomous racing, can be daunting for researchers. To help overcome this barrier, we provide our full-stack autonomous control solution. Commonly, an autonomous control stack consists of four key components: perception, localisation and mapping, planning, and control [24]. Although this pattern provides boundaries that separate out the concerns of each subsystem, it creates an all-or-nothing paradigm where all components must be present for an autonomous system to demonstrate competency. To overcome the upfront technical investment of designing, implementing and testing such a system, AARK provides the ACMPC package as a starting point for researchers. ACMPC is a full stack autonomous control solution capable of controlling vehicles around several circuits. This particular solution is an improved version of the controller that won the 2022 Learn 2 Race Challenge. A high level diagram of ACMPC is shown in Figure 3. ACMPC uses a semantic segmentation model to perceive the local track profile. These local precepts are used to map the circuit, localise the agent, and plan a trajectory using model predictive control (MPC). For a detailed explanation of the control solution see our previous work: The Edge of Disaster: A Battle Between Autonomous Racing and Safety [25]. The specific implementation contained in ACMPC focuses on the independent and modular nature of each component in the autonomous control stack. This allows researchers to focus their efforts on a single piece of the system. The entire stack is written in Python and each component is contained in a separate process to allow performant parallel execution. A diagnostic dashboard which streams visualisations and status information while agents are driving is included to aid debugging and system analysis. A Twitch stream of ACMPC racing on a rotation of tracks can be found at https://www.twitch.tv/vid3ojames. This stream is a live real-time feed running on a desktop computer equipped with a Ryzen 5950X and RTX 4090.
@@ VI-B Evaluation
[…]
@@
VI-B1 Single Lap Speed
@@
Agents are permitted any number of warm-up laps where they are able to bring the vehicle up to operational temperature and get a flying start. This enables an iterative improvement from agents that start from a standstill to those that use practice time to perform mapping, leverage the increased grip gained from managing their tyres, and carry speed from a previous lap into their hotlap.
VI-B2 Five Lap Consistency
@@
By extending the evaluation window from one to five consecutive laps, and averaging the achieved lap times, a better view of how well the agent can maintain race pace and its consistency as a driver can be discerned.
VI-B3 Twenty Lap Conservation
@@
Expanding the number of laps that are averaged across, to a common length between pit-stops, further increases complexity. Over this time period, agents that learn to manage tyre temperatures, pressures, and degradation while leveraging the decrease in platform mass due to fuel consumption, will outperform more naive solutions.
VII LIMITATIONS AND FUTURE WORK
@@
In it’s current state, AARK supports one game, AC. However, AARK’s implementation is as simulator agnostic as possible. Many popular racing games, such as Assetto Corsa Competizione and iRacing, expose real-time vehicle state information for use in motion platforms, steering feedback, ABS/TC vibrations, and custom dashboards. Although ACDG does rely on AC specific features, an agent trained on ACDG generated data can be tested on other simulators by adapting ACI such that frames are captured from a different application window and state information is collected accordingly. Therefore, as new racing simulators are released, autonomous control solutions developed using AARK can continue to be evaluated and compared on new simulators into the future.
@@
Support for multiple autonomous agents competing against one another in online lobbies, ROS integration, longer format races with pit stops, more circuits, and different motorsport formats such as autocross and rally are all on AARK’s development road map. Enabling real-time emulation of LiDAR and production of ground truth machine vision annotations by moving ray casting operations to a co-processor is another key milestone. These additional features will further expand the types of research that can be done using AARK.
VIII CONCLUSION
@@
AARK lays a foundation for autonomous racing research that presents many exciting and diverse opportunities. In our research, AARK provides latitude to explore advanced mapping, localisation and control techniques that require increased simulator fidelity. AARK provides an accessible, realistic, and customisable autonomous racing agent development kit bridging the gap between RL based and classical system design philosophies. By providing an open-sourced modular full stack autonomous system in the form of ACMPC, AARK’s accessibility transcends price and usability, researchers can get started immediately by branching off from a working system. AC’s realism and extensive modification community make it the perfect starting simulator, ensuring adaptability to accommodate many different types of approaches to autonomous racing systems. Finally, AARK is designed as a generic simulation interface that makes swapping out components, such as AC for another simulator, possible. Ultimately, AARK allows researchers to get closer, sooner, to demonstrating capable software autonomous system to explore algorithmic approaches and justify investment into their autonomous programs prior to the acquisition of physical platforms or proprietary commercial simulators.
@@ VIII-A Implementation details of ACI
To allow an autonomous agent to interact with AC, ACI provides bundled captured frames with state information from the simulator to autonomous systems and allows those systems to input vehicle controls in response. As AARK has a strong emphasis on accessibility it was important to ensure execution of simulation and autonomous systems are as flexible as possible. For example, if a researcher prefers Linux operating systems ACI enables the programmatic execution and configuration of AC on Linux machines via the use of compatibility tools. So far support for using Cross Over and Steam Proton are included in ACI. These compatibility tools create a separate environment for execution of the windows native applications and therefore cannot be directly accessed by the Linux host. This air gap between the host and the simulator is bridged by ACI’s state server, which packages raw state information and forwards it to a client on the host via a socket. In this configuration where both simulator and autonomous system are co-located on a host machine, an extremely low latency between observations, processing and control is achieved. Equally, this approach to implementation allows for a network to be between the simulator’s host and the autonomous system’s; enabling exploration of hardware in the loop or distributed system based approaches. Further to this, ACI is provides bindings which allow an autonomous system to be executed inside a docker container, vastly improving the development experience and ease of reproducibility. Figure 5, shows a detailed overview of the main components of ACI and ACDG indicating their relationship to one another. For specific details and usage visit AARK’s documentation website here: https://www.adelaideautonomous.racing/docs-aci/
VIII-B Included for convenience
@@
Included with AARK are several quality of life enhancements that have made development and tuning of our autonomous racing agent more intuitive and convenient. These internal tools are provided to the community to help researchers get started in such a complex field of research. ACMPC includes a standalone PyQt diagnostics dashboard/system monitor, a non-interactive version of which can be seen on the left of Figure 1. This dashboard is capable of visualising system behaviours such as localisation, segmentation, and planning. Visualisations are computed in a separate processes from the autonomous system itself allowing for non-blocking execution of critical control processes; minimising visualisations impact on system performance. In a separate repository code for training and evaluation semantic segmentation models on datasets produced by ACDG is provided. This allows researchers to reproduce reported results from scratch, improve on or experiment with approaches to modelling and train on new data generated for novel circuits. An additional telemetry format is supported by ACI to allow convenient storage and access by data scientists. ACI does this by providing a telemetry writer that stores simulation state information in a PostgreSQL database. This allows interaction and manipulation of telemetry data via SQL which has been useful for creating custom analysis metrics or windows for those less familiar with MoTeC. Lastly, a package that is able to read and assemble extracted simulation vehicle data into a python object is included. This provides programmatic access to vehicle information such as wheelbase and suspension geometry for all simulation included and user created vehicles in AC. On top of this reader a series of mathematical convenience functions are provided to translate steering inputs into tyre angles.

### v2

@@ V-A2 Vision
The datasets used to train ACMPC’s perception model are outlined in Table I. These provide semantic, depth and normal map annotations for each of AARK’s supported racetracks. Semantics consist of nine race track specific classes. Seven of these classes describe the type of surface: road, drivable, curb, carpet, grass, sand and water. Drivable is reserved to delineate the portion of road surface that is legal for vehicles to traverse while racing. A track limits label is included that allows models to perceive road markings which can indicate different features of importance around a racecourse. The remaining classes denote physical objects in the environment: people, vehicles, structures and vegetation. Structures encompass a super set of buildings, fences, safety barriers and signs. ACDG is capable of generating ground truth annotations for the following machine vision tasks: semantic segmentation, depth, normal maps, semantic LiDAR, and 3D object detection. Currently AARK provides pre-generated datasets for the first three of these tasks.
@@ VI-A ACMPC
Autonomous racing research carries with it a significant barrier to participation that can be daunting for researchers. On such barrier is the all-or-nothing nature of control solutions, requiring large amounts of upfront platform development before being able to experiment. To help overcome this, AARK includes a full-stack autonomous control solution. Commonly, an autonomous control stack consists of four key components: perception, localisation and mapping, planning, and control [24]. Although this pattern provides boundaries that separate out the concerns of each subsystem, it creates an all-or-nothing paradigm where all components must be present for an autonomous system to demonstrate competency. To overcome the upfront technical investment of designing, implementing and testing such a system, AARK provides the ACMPC package as a starting point for researchers. ACMPC is a full stack autonomous control solution capable of controlling vehicles around several circuits. This particular solution is an improved version of the controller that won the 2022 Learn 2 Race Challenge. A high level diagram of ACMPC is shown in Figure 3. ACMPC uses a semantic segmentation model to perceive the local track profile. These local precepts are used to map the circuit, localise the agent, and plan a trajectory using model predictive control (MPC). For a detailed explanation of the control solution, see our previous work: The Edge of Disaster: A Battle Between Autonomous Racing and Safety [25]. The specific implementation contained in ACMPC focuses on the independent and modular nature of each component in the autonomous control stack. This allows researchers to focus their efforts on a single piece of the system. The entire stack is written in Python and each component is contained in a separate process to allow performant parallel execution. A diagnostic dashboard which streams visualisations and status information while agents are driving is included to aid debugging and system analysis. A Twitch stream of ACMPC racing on a rotation of tracks can be found at https://www.twitch.tv/vid3ojames. This stream is a live real-time feed running on a desktop computer equipped with a Ryzen 5950X and RTX 4090.
@@ VI-B Evaluation
TABLE II: Lap times achieved by ACMPC agents at different tracks in an Audi R8 LMS 2016. Times are expressed as m:s.ms $\pm$ s.ms. Results for multi-lap benchmarks show the average lap time plus minus one standard deviation.
@@
Single Lap Speed: Agents are permitted any number of warm-up laps where they are able to bring the vehicle up to operational temperature and get a flying start.
@@
This enables improvement over agents that start from a standstill to those that use practice time to perform mapping, leverage the increased grip gained from managing their tyres, and carry speed from a previous lap into their hotlap.
@@
Five Lap Consistency: By extending the evaluation window from one to five consecutive laps, and averaging the achieved lap times, a better view of how well the agent can maintain race pace and its consistency as a driver can be discerned.
@@
Twenty Lap Conservation: Expanding the number of laps that are averaged, to a common length between pit-stops, further increases complexity. Over this time period, agents that learn to manage tyre temperatures, pressures, and degradation, while leveraging the decrease in platform mass due to fuel consumption, will outperform more naive solutions.
TABLE III: Lap times achieved by SAC agents at different tracks in an Audi R8 LMS 2016. Times are expressed as m:s.ms $\pm$ s.ms. Results for multi-lap benchmarks are not shown due to the SAC agents inability to achieve the required number of consecutive laps.
Monza Spa Silverstone Vallelunga Bathurst Imola
1:58.072 2:31.457 2:23.979 1:38.765 2:12.865 1:53.304
VII Reinforcement Learning
VII-A Soft Actor Critic
To demonstrate ACI’s compatibility with reinforcement learning based control solutions, a gym interface and modified version of the Soft Actor Critic (SAC) agent implementation in  [15] is provided as part of AARK. Although SAC may seem to achieve a better level of single lap control than ACMPC, there are significant differences in experimental conditions which prevent an apples to apples comparison of the two techniques. Most glaringly, the SAC agent, at both training and inference, receives a large amount of ground truth information from the simulation. This includes information such as the curvature of a reference racing line for the next 300 meters ahead of the vehicle’s current position, the amount of slip each tyre is experiencing against the road surface, ego angular and axial velocities, and a collection of distance measurements from the vehicle’s current position to the reference racing line and track limits. In contrast ACMPC executes vehicle control based on the current speed of the vehicle and a camera image. Experimentation on SAC agents under evaluation conditions were unable to produce the same level of consistency as ACMPC. SAC agents struggle to produce five consecutive laps without exceeding tracklimits or crashing, let alone the target duration of twenty. AARK’s version of the SAC agent provided by [15] makes significant optimisations to the procedures used for calculating state representations and training stability. The state calculating subroutines no longer requires a GPU to run in real time, significantly increasing the accessibility of the approach to those with less computational resourcing. Additionally, it was found that training stability was greatly improved by layer a curriculum based approach on top of SAC. The curriculum encourages agents to learn the racing line and complete laps before attempting to travel faster. This is done by limiting the reward received during training for going faster until an agent has completed a replay buffer’s worth of successful episodes. From this point, reward for pace is increased each time a successful episode is completed. This process is repeated until reward derived from speed is effectively uncapped. Without this warm-up agent become stuck in a local reward maxima, where they fail to link a full lap together. Usually this manifests as a failure to learn to brake into tight corners, with agents greedily prioritising speed over positioning and longevity. A series of single lap baseline speeds for SAC agents trained on various circuits are reported in Table III.
VIII LIMITATIONS AND FUTURE WORK
@@
In its current state, AARK supports one game, AC. However, AARK’s implementation is as simulator agnostic as possible. Many popular racing games, such as Assetto Corsa Competizione and iRacing, expose real-time vehicle state information for use in motion platforms, steering feedback, ABS/TC vibrations, and custom dashboards. Although ACDG relies on AC specific features, an agent trained on ACDG generated data can be tested on other simulators by adapting ACI such that frames are captured from a different application window and state information is collected accordingly. Therefore, as new racing simulators are released, autonomous control solutions developed using AARK can continue to be evaluated and compared on new simulators into the future.
@@
Support for multiple autonomous agents competing against one another in online lobbies, ROS integration, longer format races with pit stops, more circuits, and different motorsport formats such as autocross and rally are all on AARK’s development road map. Enabling real-time emulation of LiDAR and production of ground truth machine vision annotations, by moving ray casting operations to a co-processor, is another key milestone. These additional features will further expand the types of research that can be done using AARK.
Finally, the simulation to real gap that exists between any digital twin and reality has not been fully evaluated in the case of Assetto Corsa and reality. In the future it would be possible to explore this explicitly if provided with access to an autonomous racing platform and circuit.
IX CONCLUSION
@@
AARK lays a foundation for autonomous racing research that presents many exciting and diverse opportunities. In our research, AARK provides us the latitude to explore advanced mapping, localisation, and control techniques that require increased simulator fidelity. For the wider community, AARK provides an accessible, realistic, and customisable autonomous racing agent development kit bridging the gap between reinforcement learning and classical system design philosophies. By providing an open-source, modular, full stack autonomous system in the form of ACMPC, researchers can get started immediately by branching off from a working system. AC’s realism and extensive modification community make it the perfect starting simulator, ensuring adaptability to accommodate many different types of approaches to autonomous racing systems. Finally, AARK is designed as a generic simulation interface that makes swapping out components, such as AC for another simulator, possible. Ultimately, AARK allows researchers to get closer, sooner, to demonstrating capable autonomous system to explore algorithmic approaches and justify investment into their autonomous programs prior to the acquisition of physical platforms or proprietary commercial simulators.
@@ IX-A Implementation details of ACI
To allow an autonomous agent to interact with AC, ACI provides bundled captured frames with state information from the simulator to autonomous systems and allows those systems to input vehicle controls in response. As AARK has a strong emphasis on accessibility it was important to ensure execution of simulation and autonomous systems are as flexible as possible. For example, if a researcher prefers Linux operating systems ACI enables the programmatic execution and configuration of AC on Linux machines via the use of compatibility tools. So far, support for using Cross Over and Steam Proton are included in ACI. These compatibility tools create a separate environment for execution of the Windows native applications and therefore cannot be directly accessed by the Linux host. This air gap between the host and the simulator is bridged by ACI’s state server, which packages raw state information and forwards it to a client on the host via a socket. In this configuration, where both simulator and autonomous system are co-located on a host machine, an extremely low latency between observations, processing and control is achieved. Equally, this approach to implementation allows for a network connection between the simulator and autonomous system; enabling exploration of hardware in the loop or distributed system based approaches. Furthermore, ACI provides bindings which allow an autonomous system to be executed inside a docker container, vastly improving the development experience and ease of reproducibility. Figure 5, shows a detailed overview of the main components of ACI and ACDG indicating their relationship to one another. For specific details and usage visit AARK’s documentation website here: https://www.adelaideautonomous.racing/docs-aci/
IX-B Included for convenience
@@
Included with AARK are several quality of life enhancements that have made development and tuning of our autonomous racing agent more intuitive and convenient. These internal tools are provided to the community to help researchers get started in such a complex field. ACMPC includes a standalone PyQt diagnostics dashboard/system monitor, a non-interactive version of which can be seen on the left of Figure 1. This dashboard is capable of visualising system behaviours such as localisation, segmentation, and planning. Visualisations are computed in a separate processes from the autonomous system itself, allowing for non-blocking execution of critical control processes; minimising visualisations impact on system performance. In a separate repository, code for training and evaluation semantic segmentation models on datasets produced by ACDG is provided. This allows researchers to reproduce reported results from scratch, improve on or experiment with approaches to modeling and train on new data generated for novel circuits. An additional telemetry format is supported by ACI to allow convenient storage and access by data scientists. ACI does this by providing a telemetry writer that stores simulation state information in a PostgreSQL database. This allows interaction and manipulation of telemetry data via SQL which has been useful for creating custom analysis metrics or windows for those less familiar with MoTeC. Lastly, AC-Extras (ACE) is able to read and assemble extracted simulation vehicle data into Python objects is included. This provides programmatic access to vehicle information such as wheelbase and suspension geometry for all simulated vehicles in AC. ACE is also packaged with a series of mathematical convenience functions which use parsed vehicle geometry information to calculate transformations between steering inputs and tyre angles.